OpenStudy (superhelp101):

When an alpha particle is released by an atom, the atom has an increased mass number of four. gains a neutron. loses one electron. loses a helium nucleus.

OpenStudy (superhelp101):

@Abhisar

OpenStudy (superhelp101):

Idk much about...

OpenStudy (superhelp101):

this

OpenStudy (abhisar):

An alpha particle is nothing but 2 Protons+2Neutrons \(\huge \therefore\) Alpha Particle = Helium Nucleus

OpenStudy (superhelp101):

if it has to 2Neutrons then it gains nertrons is right ??

OpenStudy (abhisar):

If anything emits/releases alpha particle then it loses neutrons and protons

OpenStudy (superhelp101):

oh hmm.. but the choices for "losing" are only loses one electron and loses a helium nucleus sorry i confused

OpenStudy (abhisar):

Yes...so loses a helium nucleus will be the correct option

OpenStudy (abhisar):

because alpha particle is equivalent to a helium nucleus which also contains 2 neutrons and 2 protons

OpenStudy (abhisar):

Got it ?

OpenStudy (superhelp101):

Yeps!! I got it now Thxxx!

OpenStudy (abhisar):

\(\rlap{\color{red}{\huge\bigstar}\huge \color{red}{ \text{You're Most Welcome! }}\color{red}\bigstar}{\; \color{black}{\huge\bigstar}\huge \color{black}{\text{You're Most Welcome! }}\color{black}\bigstar}\) \(~~~~~~~~~~~~~~~~~~~~~~~~~~~\color{green}{\huge\ddot\smile}\color{blue}{\huge\ddot\smile}\color{pink}{\huge\ddot\smile}\color{red}{\huge\ddot\smile}\color{yellow}{\huge\ddot\smile}\)

OpenStudy (superhelp101):

oooo :o the new and advanced "You're Most Welcome" pretty :3

OpenStudy (abhisar):

\(\Huge{\overset{\frown}{\normalsize \left( \begin{matrix} \Large\cdot \quad \cdot\\ \cdot\\ \huge \smile \end{matrix} \right)}}\normalsize \\ \;/\quad \;\;\quad \backslash\)